OpenStudy (anonymous):

Prove F(x)=sin^2(x) and G(x)=cos^2(x) have the same derivatives. What can you say about the difference f-g? Explain using a consequence of the Mean Value theorem?

OpenStudy (anonymous):

whoa

OpenStudy (irishboy123):

F + G = 1 F' + G' = 0 F' = - G' **same** derivatives??

OpenStudy (anonymous):

Another way of showing the claim is wrong: \[\sin^2x=\frac{1-\cos2x}{2}~~\implies~~\frac{d}{dx}\sin^2x=\sin2x\\ \cos^2x=\frac{1+\cos2x}{2}~~\implies~~\frac{d}{dx}\cos^2x=-\sin2x\]